OpenStudy (anonymous):

A single die is rolled. Find the probability of rolling an even number or a number less than 6. --Is the Probability 1/3, or something else?

OpenStudy (anonymous):

A number less than six is pretty much gonna happen all the time unless you land on six.

OpenStudy (anonymous):

An even number is 50%.

OpenStudy (anonymous):

Did this help?

OpenStudy (anonymous):

Still a bit confused on this one, Would it be like a 5/6 + 1/2 = 8/6 = 1 1/6 chance of happening?

OpenStudy (anonymous):

I have a feeling it is 1/3, but I'm not sure.

OpenStudy (anonymous):

Can anyone confirm this?

OpenStudy (anonymous):

an even number would've been 1/2 and a number less than six is 5/6? Is that what you were asking?